# UNGER: Generative Recommendation with A Unified Code via Semantic and Collaborative Integration

### v2

@@ 5. CONCLUSION AND FUTURE WORK
In this paper, we introduce a novel framework, UNGER, aimed at leveraging a unified code to encode knowledge from both semantic and collaborative modalities in generative recommendation systems. UNGER is composed of three key components: (1) A unified code generation framework that integrates semantic and collaborative knowledge, addressing the semantic domination issue by effectively capturing their complementary characteristics. (2) A joint optimization of cross-modality knowledge alignment and next item prediction tasks and a learnable modality adaption layer with AdaLN, which integrate collaborative and semantic knowledge into integrated embeddings, enabling the unified code to represent both modalities information cohesively. (3) An intra-modality knowledge distillation task with a specially designed token, which compensates for the information loss caused by the quantization process and further improves auto-regressive generation quality. Extensive experiments compared with state-of-the-art methods, combined with detailed analyses, validate the effectiveness and robustness of UNGER. In future work, we plan to leverage more available features from the semantic modality, such as item visual features, to further enhance the model’s generative recommendation performance.

### v3

@@ 5. CONCLUSION AND FUTURE WORK
In this paper, we introduce a novel framework, UNGER, aimed at leveraging a unified code to encode knowledge from both semantic and collaborative modalities in generative recommendation systems. UNGER is composed of three key components: (1) A unified code generation framework that integrates semantic and collaborative knowledge, addressing the semantic domination issue by effectively capturing their complementary characteristics. (2) A joint optimization of cross-modality knowledge alignment and next item prediction tasks and a learnable modality adaption layer with AdaLN, which integrate collaborative and semantic knowledge into integrated embeddings, enabling the unified code to represent both modalities information cohesively. (3) An intra-modality knowledge distillation task with a specially designed token, which compensates for the information loss caused by the quantization process and further improves auto-regressive generation quality. Extensive experiments compared with state-of-the-art methods, combined with detailed analyses, validate the effectiveness and robustness of UNGER.
In future work, we plan to extend UNGER along two complementary directions. The first direction involves enriching the semantic modality by incorporating visual embeddings extracted from item images or video frames using pretrained deep neural networks (He et al., 2016; Radford et al., 2021) and Vision Transformers (Dosovitskiy et al., 2020), as well as audio and speech representations from domains such as music and podcasts. This enhancement will enable the unified code to capture stylistic, aesthetic, and affective aspects of user preferences that are not fully conveyed by textual descriptions alone. The second direction focuses on improving the tokenizer by adopting advanced vector quantization techniques from other fields, including LFQ (Yu et al., 2023), IBQ (Shi et al., 2024), and ActionPiece (Hou et al., 2025), which provide better convergence properties, higher codebook utilization, and more expressive latent representations. These improvements are expected to further enhance the generation quality and robustness of the model.
It is also worth noting that the current evaluation is restricted to relatively sparse subsets of the Amazon dataset. While this choice is consistent with prior work, it does imply that the behavior of UNGER on denser recommendation scenarios has not yet been systematically explored. We leave such evaluations as an important avenue for future investigation, which will help establish the broader applicability and generalizability of the proposed framework.
Acknowledgements.
This work is supported by the National Key Research and Development Program of China under grant 2024YFC3307900; the National Natural Science Foundation of China under grants 62302184, 62376103, 62436003 and 62206102; Major Science and Technology Project of Hubei Province under grant 2024BAA008; Hubei Science and Technology Talent Service Project under grant 2024DJC078; Ant Group through CCF-Ant Research Fund. The computation is completed in the HPC Platform of Huazhong University of Science and Technology.